equation of a parabola with a focus of (0,3) and a directrix of y=-3
Answer by Alan3354(69443)

equation of a parabola with a focus of (0,3) and a directrix of y=-3
-------------
All points on the parabola are equidistant from (0,3) the y = -3
----
For point (x,y)
d = sqrt((y-3)^2 + x^2) = y+3
Square both sides
(y-3)^2 + x^2 = y^2 + 6y + 9
y^2-6y+9 + x^2 = y^2+6y+9
x^2 = 12y
y = x^2/12